# 3D Voxel Maps to 2D Occupancy Maps for Efficient Path Planning for Aerial and Ground Robots

### v1

@@ III-C2 UGV map $\mathcal{M}_{G}$
The 2D map generation for ground-based robots utilizes the same free space and occupancy detection as the map generated for the UAV, but in addition, the slope map generated in subsection III-B is used as well. In the UGV map $\mathcal{M}_{G}$, if the slope associated with a cell that is free is greater than $R_{MS}$, the cell is instead considered occupied, where $R_{MS}$ is the maximum slope a ground robot can climb. Thus, the low walls and scattered objects that are relevant for UGVs are detected and represented in the UGV map $\mathcal{M}_{G}$.
IV Use Cases and Utility of the Method
The authors identify two distinct use cases for the 3D to 2D map conversion method proposed in this work. The first is to minimize computations and memory usage for both UAVs and UGVs in large-scale navigation scenarios, and the second is to reduce the data transmitted between multiple robotic agents in bandwidth-limited scenarios.
In the first case, the 2D maps can be used to find 3D paths for UAVs and UGvs for efficient large-scale navigation. A general path is initially determined for the robots on the 2D map, using any of the classical or state-of-the-art occupancy-grid-based path planning methods, followed by elevation of the path into 3D space through the approach described in section V. Path planning using the 2D map is more efficient with respect to memory and CPU usage. This, as a 2D map, is significantly lighter on memory, and navigation problems are quicker to solve, as the $Z$-axis need not be considered. For very precise local navigation, a local section of the voxel map in a small neighborhood of the robot can be used, as there is no necessity to load the complete voxel map. This strategy also allows the robot to maintain a higher-resolution local voxel map, as it only needs to store a small portion of the complete map in its memory.
The second case addresses scenarios where multiple robotic agents need to communicate and exchange local sections of a larger map with each other. Since the 2D map generally has a smaller file size, it can be transmitted between robots more rapidly within the allocated bandwidth. Although the 2D map, along with the height map, does not perfectly represent the environment, it still offers a good practical sense of the layout, enabling the robots to perform general path planning, as stated previously.
V 2D to 3D Path Conversion
Figure 3: A sphere-shaped safety region is used to find collision-free paths for the UAVs.
As discussed in Section IV, one of the primary advantages of the proposed method is its ability to plan paths using a 2D map instead of relying on a full 3D voxel map. Let us assume that a classical or state-of-the-art 2D path planner is used to find paths in the UAV and UGV maps. To make the 2D paths practicable for robots, especially UAVs, it is essential to determine the height, the $Z$ position, of each point along the path. To address this challenge, the following process is employed to determine the $Z$ position of each point on a 2D path $\mathcal{P}$ generated on the 2D occupancy maps.
@@ IV Use Cases and Utility of the Method
where ˇ$\check{\mathcal{H}}(\mathcal{P}_{i})$ represents the floor height in the height map $\mathcal{H}$ at the position of $\mathcal{P}_{i}$. The parameter $p_{f}$ is user-defined, specifying the number of steps that are considered in front of and behind the robot along the path. The $R_{off}$ is another user-defined parameter that sets a desired safe height above the ground for the path. The use of the $\max$ function ensures that the path adjusts its height $p_{f}$ steps before encountering an obstacle. For generating 3D paths for the UAVs, an additional check is performed to avoid collisions with the safety region around the robot. The method uses a sphere-shaped safety region for UAVs, as illustrated in Figure 3. The safety region has a radius of $0<R_{r}\leq 2*R_{maxZ}$. If the height calculated in the previous step results in a collision with either the floor or the ceiling in the height map $\mathcal{H}$, the point in the path that results in a collision is moved to a $Z$ position that will not cause a collision, ignoring $R_{off}$ at that point in the path.
@@ VI Validation
The proposed method was validated in two environments: a cave environment and an indoor environment. A voxel map was generated with a resolution of $\mathcal{V}_{res}=0.1$ m in the cave environment and with a resolution of $\mathcal{V}_{res}=0.05$ in the indoor environment. Using the map conversion method outlined in Section III, 2D occupancy maps $\mathcal{M}_{A}$ and $\mathcal{M}_{G}$ were generated from voxel maps of the two environments. The 2D maps for the Unmanned Aerial Vehicle (UAV) are shown in Figures 4(a) and 5(a), and for the Unmanned Ground Vehicle (UGV) are shown in Figures 4(b) and 5(b). Note that low walls and scattered objects are not relevant for aerial robots, as there is navigable space for aerial robots above the walls and scattered objects. For the ground robots, however, the walls and scattered objects are highly relevant and are retained by the proposed method, using the information contained in the slope map (the procedure is described in Section III-C).
The height map $\mathcal{H}$, specifically $\check{h}_{x_{i},y_{i}}$, is utilized to find the floor and the ceiling of the map. The height of the floor, with respect to the origin of the voxel map, is depicted in Figures 4(c) and 5(c). Using the height maps, slope estimation described in Section III-B is used to find the slope in the neighborhood of every grid cell. The slope maps are depicted in Figures 4(d) and 5(d). The slope maps, along with the UAV maps, are used to find the 2D occupancy maps for the ground robots, which are shown in Figures 4(b) and 5(b). In the slope map, the boundary of low walls and obstacles can clearly be seen. In the UAV maps, the slope thresholding approach using the parameter $R_{MS}$ (described in Section III-C) is used to assign the cells associated with the boundaries of walls and scattered objects an occupancy value of 1. Two 2D paths were generated using a 2D planning algorithm: one for the Unmanned Aerial Vehicle (UAV), as shown in Figures 4(a) and 5(a), and another for the Unmanned Ground Vehicle (UGV), as seen in Figures 4(b) and 5(b). The 3D paths for both robots are generated using the method in V, as shown in the figures 4(e), 4(f), 5(e), and 5(f). The parameters used in 3D to 2D map conversion and for 3D path generation are listed in Table I.
@@
$R_{maxZ}$ $O_{max}$ $R_{MS}$ $R_{r}$ $R_{off}$ $p_{f}$
UAV 1 m 0.1 NA 0.5 m 1 m 1 m / $\mathcal{V}_{res}$
UGV 1 m 0.1 0.35 NA 0.1 m 1 m / $\mathcal{V}_{res}$
@@ VII Discussion
As demonstrated in Section VI, the method effectively converts voxel maps in environments featuring non-uniform height variations and in the presence of overhangs/ceilings. The UGV map includes obstacles and walls impassable by the UGV, thus ensuring comprehensive environmental representation. However, a limitation of the traversability slope estimation discussed in Section III-B is its dependency on having a sufficiently low resolution in the voxel map. In the validation of the proposed method, a resolution of $\mathcal{V}_{res}=0.1$ m and $\mathcal{V}_{res}=0.05$ was utilized, which is relatively fine for voxel mapping. Nonetheless, as argued in Section IV, by using a 2D map for global navigation, only a small local portion of the 3D map is needed for local navigation. This allows the local map to have a higher resolution without significant memory cost compared to needing to have the full voxel map loaded.
A further issue with the UGV map is the inaccurate representation at some frontiers, which form the boundary between free and unknown space. This inaccuracy arises from the 3D mapping solutions, which do not map the floor correctly at the edges of the Lidar. Thus, the greedy method for free-space detection presented in Section III-A results in a sudden height increase in the area near the frontiers. Despite this limitation, the mapped area remains accurate for UGV navigation. In scenarios such as autonomous exploration, reliable frontier detection is an absolute necessity. The limitation implies that the UGV-generated map is unsuitable in the exploration case, thus necessitating reliance on the local voxel map for frontier detection.
One limitation of using a 2D planner for UAV path planning instead of a full 3D planner is that it overlooks height variations in the environment. While this may not be problematic in general, this limitation can lead to suboptimal path choices, as circumventing a tall obstacle might be more efficient than flying over it. A potential solution could be to integrate height differences into the 2D planner’s calculations, thus utilizing the height map produced by the method more effectively, and this would form one of our future research directions.
@@ VIII Conclusions
Motivated by the expensive cost of UAV and UGV navigation planning over large-scale 3D maps, both in terms of CPU and memory, this work presented a novel method for converting 3D voxel maps to 2D occupancy maps for ground and aerial robots, along with an accompanying height map. By allowing the robots to use a 2D map for large-scale navigation, CPU usage can be minimized by using computationally efficient 2D planners instead of full 3D planners. This, combined with the fact that a 2D map uses less memory than a 3D map, results in a more computationally efficient large-scale navigation stack for an autonomous robot. A method was also proposed for converting the paths generated by the 2D planer back into 3D space using the height information extracted by the method. The method proposed in this paper is capable of generating 3D paths for both ground and aerial robots and was validated successfully in two different environments (a cave and an indoor scenario).

### v2

@@ III-C2 UGV map $\mathcal{M}_{G}$
The 2D map generation for ground-based robots utilizes the same free space and occupancy detection as the map generated for the UAV, but in addition, the slope map generated in subsection III-B is used as well. Cells near low walls and scattered objects exhibit increased slope values as the height increases suddenly. In the UGV map $\mathcal{M}_{G}$, if the slope associated with a cell that is free is greater than $R_{MS}$, the cell is instead considered occupied, where $R_{MS}$ is the maximum slope that a ground robot can traverse. Thus, the low walls and scattered objects that are relevant for UGVs are detected and represented in the UGV map $\mathcal{M}_{G}$.
IV 2D to 3D Path Conversion
As discussed in Section I, one of the primary advantages of the proposed method is its ability to plan paths using a 2D map instead of relying on a full 3D voxel map. Let us assume that a classical or state-of-the-art 2D path planner is used to find paths in the UAV and UGV maps. To make the 2D paths practicable for robots, especially UAVs, it is essential to determine the height, the $Z$ position, of each point along the 2D path. To address this challenge, the following process is employed to determine the $Z$ position of each point on a 2D path $\mathcal{P}$ generated on the 2D occupancy maps.
@@
where ˇ$\check{\mathcal{H}}(\mathcal{P}_{i})$ represents the floor height in the height map $\mathcal{H}$ at the position of $\mathcal{P}_{i}$. The parameter $p_{f}$ is user-defined, specifying the number of steps that are considered along the path in front of and behind the current position of the robot. The parameter $R_{off}$ is another user-defined parameter that sets a desired safe height above the ground for the path. The use of the $\max$ function ensures that the path proactively adjusts its height $p_{f}$ steps before encountering an obstacle. For UAVs, considering just the points along the path for height computation may prove insufficient to ensure safe distances from the objects in the neighborhood of the path. Therefore, when generating 3D paths for the UAVs, an additional check is performed to avoid collisions with a sphere-shaped safety region around the UAVs. The safety region has a radius of $20<R_{r}\leq R_{maxZ}/$ 2. If the height calculated in the previous step results in a collision with either the floor or the ceiling in the height map $\mathcal{H}$, the point in the path that results in a collision is moved to a $Z$ position that does not cause a collision, ignoring $R_{off}$ at that point in the path.
V Use Cases and Utility of the Proposed Method
The authors identify several use cases for the 3D to 2D map conversion method proposed in this work. To begin with, a 2D map on a computer screen can be more intuitive for a human compared to a 3D map, to both visualize and interact with, for example, in selecting goal points or monitoring the navigation of a robot. The second use case arises in using a 2D map for global navigation for UAVs. Path planning using the 2D map is more efficient with respect to memory and CPU usage, as the navigation problems are quicker to solve since the $Z$-axis need not be considered. The third utility of the proposed methods is in enabling the use of methods (such as map segmentation or several 2D navigation solutions) that are designed specifically for 2D maps or scale poorly for a 3D map, to be applicable in 3D scenarios.
Map regions, objects, and paths found using the 2D maps can then be moved to 3D space using the height information of the environment extracted by the proposed method. This can be done using the method presented in Section IV, which converts 2D paths planned on a 2D map to 3D paths for both UGVs and UAVs.
The final utility arises in scenarios where multiple robotic agents need to communicate and exchange large maps with each other. Since the 2D map has a smaller file size than the full 3D Voxel map (as shown in Table II in Section VI), it can be communicated between robots or to a remote location more rapidly within the allocated bandwidth. Although the 2D map and height map do not perfectly represent the environment, they still offer a good practical sense of the layout, enabling the robots to perform general path planning.
@@ VI Validation
The proposed method was validated using two 3D voxel maps of real environments gathered from robotic experiments: a small section of a cave and a large outdoor environment. A voxel map using the UFOMaps [15] was generated with a resolution of $\mathcal{V}_{res}=0.1$ m in the cave environment and with a resolution of $\mathcal{V}_{res}=0.2$ in the outdoor environment. Using the map conversion method outlined in Section III, 2D occupancy maps $\mathcal{M}_{A}$ and $\mathcal{M}_{G}$ were generated from voxel maps of the two environments.
All scenarios were executed, and computation times were measured on the same computer. The tests were run on a single thread of an AMD 5850U CPU using Linux kernel version 6.9.7. The parameters used by the method are listed in Table I.
@@
$R_{maxZ}$ $O_{min}$ $S_{A}$ $R_{MS}$ $R_{r}$ $R_{off}$ $p_{f}$
UAV 1 m 0.5 NA 2 0.5 m 1 m 2 m / $\mathcal{V}_{res}$
UGV 1 m 0.5 0.2 2 NA 0.1 m 0.5 m / $\mathcal{V}_{res}$
VI-A Conversion of static 3D voxel maps
The 2D maps for the Unmanned Aerial Vehicle (UAV) are shown in Figures 4(a) and 5(a), and for the Unmanned Ground Vehicle (UGV) are shown in Figures 4(b) and 5(b). Note that low walls and scattered objects are not relevant for aerial robots, as there is navigable space for aerial robots above the walls and scattered objects. For the ground robots, however, the walls and scattered objects are highly relevant and are retained by the proposed method, using the information contained in the slope map (the procedure is described in Section III-C) and using the slope threshold parameter $R_{MS}$. The slope maps are depicted in Figures 4(d) and 5(d).
The slope map is generated from the height map $\mathcal{H}$. The height of the floor, $\check{\mathcal{H}}$, with respect to the origin of the voxel map, is depicted in Figures 4(c) and 5(c).
The computation time to perform the full map conversion was 0.23s for the cave environment and 6.07s for the outdoor environment.
Table II shows the reduction in the raw size of one of the 2D maps with height data compared to the original voxel map. The voxel map is stored as an octree, while the 2D map and the height data are stored in a 2D matrix.
To validate the path conversion from 2D to 3D using the 2D map $\mathcal{M}$ and the height map $\mathcal{H}$, two 2D paths were generated per environment: one for the Unmanned Aerial Vehicle (UAV), as shown in Figures 4(a) and 5(a), and another for the Unmanned Ground Vehicle (UGV), as seen in Figures 4(b) and 5(b). The paths where generated using a path-planer developed for our earlier work on exploration [16] that is designed only for a 2D map. The 2D paths are converted into 3D paths for both the robots using the method in IV. The resulting 3D paths are shown in the Figures 4(e), 4(f), 5(e), and 5(f).
TABLE II: Comparison between the raw sizes of different maps, expressed in megabytes (MB) and as a percentage compared to the original UFOMap used to generate the 2D map. The UFOMap is stored as an octree, while the 2D map and height data are stored in a 2D matrix.
Scenario UFOMap (MB) 2D Map (MB) 2D Map + Height Data (Floor) (MB) 2D Map + Height data (Floor & Ceiling) (MB)
Cave 6.7 [100%] 0.3 [4.5%] 1.6 [23.9%] 2.9 [43.3%]
Outdoor 169.7 [100%] 6.3 [3.7%] 31.7 [18.7%] 57.1 [33.6%]
VI-B Map Conversion During an Exploration Mission
In each scenario, a robot platform explores the environment, and the 2D map is updated alongside the Voxel map. During the mission, only the section of the 3D map that was updated was used for map conversion and then to update the 2D map. The computation times for the two scenarios are presented in the box plots in Figure 2. There are two times displayed in the box plot. Scenarios marked with ’*’ exclude the time spent reading the voxel map produced by UFOMap. Since this time will vary depending on the mapping framework used with the method, the time marked with ’*’ more accurately reflects the time used by the conversion method. In Figure 3, computation time relative to the total mission time of the exploration is shown.
Figure 2: The computation time to update the 2D map generated from the voxel map under two exploration missions. The scenarios marked with ’*’ are the time required by the method, excluding the time needed to read the UFOmap.
Figure 3: The computation time to update the 2D map generated from the voxel map over the total mission time, under two exploration missions. The scenarios marked with ’*’ are the time required by the method, excluding the time needed to read the UFOmap.
@@ VII Discussion
As demonstrated in Section VI, the method effectively converts voxel maps in environments featuring non-uniform height variations and in the presence of overhangs/ceilings. The UGV map includes obstacles and walls impassable by the UGV, thus ensuring comprehensive environmental representation. In existing conversion methods [6][7] for voxel-based maps, this would not have been possible, especially in outdoor scenarios where the height difference is more than 10 meters between the starting position’s ground level and the map’s lowest point. Another advantage of the proposed map conversion is in the raw file size of the generated 2D maps compared to the original voxel map, as can be seen in Table II. In cases where only UAV or UGV maps need to be transferred, the required size is less than $5\%$ of the original voxel map, which is significant in multi-agent scenarios where bandwidth is limited.
The time required to convert a static map increases with map size. The conversion time for the small cave scenario is 0.23s, whereas a larger outdoor scenario requires 6.7s. This increase is not an issue since the conversion of the full map is generally performed only once. In dynamic scenarios where the map size changes or certain areas are updated, such as during an exploration mission (presented in section VI-B), only the modified sections of the map need to be updated. In smaller enclosed environments, like a subterranean environment, the average computation time is only 0.11s, which allows for near real-time map updates, as seen in Figure 2(a). In contrast, in more open environments, such as the outdoor scenario, the conversion time increases to an average of 0.5s. Figure 3(a) illustrates that the computation time in the subterranean environment remains stable because the tunnels have approximately constant width. Conversely, in the outdoor scenario shown in Figure 3(b), the robot explores more open areas during the second half of the exploration mission, resulting in significant spikes in computation time.
An issue with the UGV map generated by the method is the inaccurate representation at some frontiers, which are the boundaries between free and unknown space. This inaccuracy arises from the greedy method for floor detection presented in Section III-A, which inaccurately classifies the floor height around the frontiers, resulting in a steep slope in these areas. Despite this limitation, the mapped area remains accurate for UGV navigation. In scenarios such as autonomous exploration, reliable frontier detection is an absolute requirement, and this necessitates reliance on the voxel map for frontier detection.
A limitation of using a 2D planner for UAV path planning instead of a full 3D planner, in scenarios with multiple 3D paths leading to a target point, is that a 2D planned may choose sub-optimal paths, such as flying over a tall obstacle instead of circumventing it. A potential solution could be integrating height differences into the 2D planner’s calculations, thus utilizing the height map produced by the method more effectively.
@@ VIII Conclusions
Motivated by the utility of converting 3D maps to 2D for global navigation, which enables the use of methods designed for 2D map in 3D environments, and faster information sharing between robots, this work presented a novel method for converting 3D voxel maps to 2D occupancy maps for UAVs and UGVs, along with accompanying height and slope maps. A method was also proposed for converting the paths generated by the 2D planer back into 3D space using the height information extracted by the method. The method proposed in this paper is capable of generating 3D paths for both ground and aerial robots and was validated successfully in two different environments (a cave and an outdoor scenario).